# Density Ratio Estimation-based Bayesian Optimization with Semi-Supervised Learning

### v4

@@ 1 Introduction
To make use of semi-supervised classifiers, we take into account two distinct scenarios with unlabeled point sampling and with a predefined fixed-size pool. For the first scenario, we randomly sample unlabeled data points from the truncated multivariate normal distribution using a minimax tilting method (Botev, 2017), in order that it is possible that a cluster assumption (Seeger, 2000) is assumed. Finally, we demonstrate that our method shows superior performance compared to the exiting methods in diverse experiments including synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and 64D minimum multi-digit MNIST search. Note that Tabular Benchmarks, NATS-Bench, and 64D minimum multi-digit MNIST search have access to fixed-size pools. To validate our methods, we provide thorough analyses on the components of DRE-BO-SSL in Section 6 and the appendices.
@@ 1.1 Over-Exploitation Problem
By the aforementioned definition, at early iterations of Bayesian optimization, a supervised classifier tends to overfit to a small size of $\mathcal{D}_{t}$ due to a relatively large model capacity. This consequence makes a Bayesian optimization algorithm highly focus on exploitation. Similar to our observation, the imbalance of exploration and exploitation in the DRE-based approaches is also discussed in the recent work (Oliveira et al., 2022; Pan et al., 2024). Moreover, the consequence mentioned above is different from the characteristics of probabilistic regression-based Bayesian optimization because the regression-based methods are capable of exploring unseen regions by dealing with uncertainties. Besides, even though a threshold ratio $\zeta$ might be able to mitigate this problem, an overconfident supervised classifier is likely to keep getting stuck in a local optimum as $y^{\dagger}$ cannot change dramatically; more detailed analysis on $\zeta$ is provided in Sections 6 and H.
@@ 2 Background and Related Work
It is a principled and efficient approach to finding a global solution of a challenging objective, e.g., expensive-to-evaluate black-box functions (Brochu et al., 2010; Garnett, 2023). To focus on a probabilistic regression model as a surrogate function, we omit the details of Bayesian optimization here; see (Brochu et al., 2010; Garnett, 2023) for details. In Bayesian optimization, Gaussian process regression (Rasmussen & Williams, 2006) is widely used as a surrogate function (Srinivas et al., 2010; Snoek et al., 2012) because of its flexibility with minimum assumptions on model and smoothness. While a Gaussian process is a probable choice, Bayesian optimization with diverse surrogates such as Student-$t$ process regression (Martinez-Cantin et al., 2018), Bayesian neural networks (Springenberg et al., 2016), and tree-based models (Hutter et al., 2011; Kim & Choi, 2022) has been proposed. An analogy between such models is that they model $p(y\mid\mathbf{x},\mathcal{D})$ explicitly, so that it can be used to define an acquisition function with the statistics of $p(y\mid\mathbf{x},\mathcal{D})$.
@@
Whereas knowing a data distribution $p(\mathbf{x})$ is important, it is difficult to directly estimate $p(\mathbf{x})$ (Sugiyama et al., 2012). For specific machine learning problems such as importance sampling (Kloek & van Dijk, 1978) and mutual information estimation (Bishop, 2006), we can bypass direct density estimation and estimate a density ratio. More recently, Rhodes et al. (2020) tackle a density-chasm problem using telescopic density-ratio estimation, which replaces an original problem with a set of logistic regression problems. Since the work (Rhodes et al., 2020) suffers from the issue on distribution shift, the recent work (Srivastava et al., 2023) proposes a density ratio estimation method with multinomial logistic regression, which is capable of mitigating the distribution shift using multi-class classification. In Bayesian optimization, Bergstra et al. (2011) have proposed a strategy with tree-structured Parzen estimators to estimate a density ratio as an alternative to probabilistic regression-based acquisition functions. In addition, the work (Tiao et al., 2021; Song et al., 2022) suggests methods with binary classifiers in order to estimate class probabilities as a density ratio; see Section 3 for details.
@@ 3 DRE-based Bayesian Optimization
where $\zeta=p(y\leq y^{\dagger})\in[0,1)$ is a threshold ratio. We need to find a maximizer of (1), by optimizing the following composite function: $h(p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})/p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t}))$, where $h(x)=(\zeta+(1-\zeta)x^{-1})^{-1}$. Since $h$ is a strictly increasing function, we can directly maximize $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})/p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t})$. In the previous work (Bergstra et al., 2011), two tree-structured Parzen estimators are used to estimate the respective densities, $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})$ and $p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t})$.
While the work (Bergstra et al., 2011) utilizes two distinct tree-structured Parzen estimators, Tiao et al. (2021) propose a method to directly estimate (1) using class-probability estimation (Qin, 1998; Sugiyama et al., 2012), which is called BORE. Since it can be formulated as a problem of binary classification in which Class 1 is a group of the top $\zeta$ of $\mathcal{D}_{t}$ and Class 0 is a group of the bottom $1-\zeta$ of $\mathcal{D}_{t}$ in terms of function evaluations, the acquisition function defined in (1) induces the following:
@@
Therefore, a class probability over $\mathbf{x}$ for Class 1 is considered as an acquisition function; it is simply denoted as the following:
@@
which is versatile for any non-negative utility function $u(y;y^{\dagger})$, where $f$ is a strictly convex function and $f^{*}$ is the convex conjugate of $f$. By the properties of LFBO, it is equivalent to an expected utility-based acquisition function. Along with the general framework, Song et al. (2022) prove that BORE is equivalent to the probability of improvement (Kushner, 1964) and LFBO with $u(y;y^{\dagger})=\mathbb{I}(y\leq y^{\dagger})$ is also equivalent to the probability of improvement, where $\mathbb{I}$ is an indicator function. Moreover, they show that their method with $u(y;y^{\dagger})=\max(y^{\dagger}-y,0)$, which can be implemented as weighted binary classification, is equivalent to the expected improvement (Jones et al., 1998).
@@ 4 DRE-based Bayesian Optimization with Semi-Supervised Learning
We adopt a multi-started local optimization technique, e.g., L-BFGS-B (Byrd et al., 1995), to solve (6). However, a flat landscape of $\pi_{\widehat{\mathbf{C}}_{t}}(\mathbf{x};\zeta,\mathcal{D}_{t},\mathbf{X}_{u})$ over $\mathbf{x}$ may occur, so that optimization performance can be degraded. To tackle this issue, a simple heuristic of randomly selecting a query point among points with identical highest class probabilities complements our method. Since the multi-started technique is utilized and the output of $\pi_{\widehat{\mathbf{C}}_{t}}$ is bounded in $[0,1]$, a flat landscape is easily recognized by comparing the outcomes of the multi-started strategy.
@@ 4.1 Label Propagation and Label Spreading
where $\beta$ is a free parameter given. As discussed in the work (Zhu & Ghahramani, 2002), we can learn $\beta$ in (8) by minimizing an entropy for propagated labels $\widehat{\mathbf{C}}$:
@@
See Figure 4 for the results of learning $\beta$ and Section E for analysis on the effects of $\beta$. By (8), we compute a transition probability from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ by $p_{ij}=w_{ij}/\sum_{k=1}^{n_{l}+n_{u}}w_{kj}$. Note that similarities $\mathbf{W}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ and transition probabilities $\mathbf{P}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ are defined, where $[\mathbf{W}]_{ij}=w_{ij}$ and $[\mathbf{P}]_{ij}=p_{ij}$. Moreover, by the definition of $p_{ij}$, $\mathbf{P}=\mathbf{D}^{-1}\mathbf{W}$, where $\mathbf{D}$ is a degree matrix whose diagonal entry is defined as the following:
@@
for $i\in[2]$, where $\mathbf{w}\in\mathbb{R}^{n_{l}+n_{u}}$ is similarities of $\mathbf{x}$ and $\mathbf{X}$ by (8). $[\widehat{\mathbf{C}}]_{:i}$ denotes $i$th column of $\widehat{\mathbf{C}}$. This inductive model is better than or equivalent to other classifiers without unlabeled data in certain circumstances; its analysis is provided in Section C.
@@ 4.2 Unlabeled Point Sampling
As described above, if unlabeled points are not available, we need to generate unlabeled points under a transductive learning scheme. However, it is a challenging problem unless we know a landscape of $\pi_{\mathbf{C}}$ adequately. Many studies (Seeger, 2000; Rigollet, 2007; Singh et al., 2008; Ben-David et al., 2008; Carmon et al., 2019; Wei et al., 2020; Zhang et al., 2022) investigate how unlabeled data can affect a model and whether unlabeled data helps improve the model or not.
@@
Assumption 4.1 (Cluster assumption in (Seeger, 2000)).
@@
where $\mathbf{l},\mathbf{u}\in\mathbb{R}^{d}$ are lower and upper bounds, $\boldsymbol{\Sigma}=\mathbf{A}\mathbf{A}^{\top}$ is a covariance matrix, $\mathbb{I}$ is an indicator function, and $Z\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}_{d})$ is a random variable. It is challenging to calculate a denominator of (15), $P(\mathbf{l}\leq\mathbf{A}^{\top}Z\leq\mathbf{u})$, and simulate from $f(\mathbf{z})$ because an integration of the denominator and an accept-reject sampling strategy from $f(\mathbf{z})$ are cumbersome in this multi-dimensional case. To effectively sample from the truncated multivariate normal distribution, we adopt the minimax tilting method (Botev, 2017). Compared to the method by Genz (1992), it yields a high acceptance rate and accurate sampling. In this paper $\boldsymbol{\Sigma}$ is set as an identity matrix, and $\mathbf{l}$ and $\mathbf{u}$ are determined by a search space. We will provide more detailed discussion on point sampling in Sections 6 and F.
@@ 5 Experiments
We compare baseline methods with DRE-BO-SSL in the following optimization problems: synthetic benchmarks for a scenario with unlabeled point sampling, and synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and minimum multi-digit MNIST search for a scenario with a fixed-size pool. Note that Tabular Benchmarks, NATS-Bench, and minimum multi-digit MNIST search are defined with a fixed number of possible solution candidates, which implies that they are considered as combinatorial optimization problems. By following the previous work (Tiao et al., 2021; Song et al., 2022), we set a threshold ratio as $\zeta=0.33$ for all experiments; the effects of a threshold ratio $\zeta$ are analyzed in Sections 6 and H. To solve (6), we use L-BFGS-B (Byrd et al., 1995) with 1,000 different initializations. All experiments are repeated 20 times with 20 fixed random seeds, where 5 initial points are given to each experiment. The sample mean and the standard error of the sample mean are reported. Other missing details including the details of the competitors of our methods are presented in Section D.
@@
  • Gaussian process, EI and UCB: It is a Bayesian optimization strategy, which is defined with Gaussian process regression with the Matérn 5/2 kernel (Rasmussen & Williams, 2006), where expected improvement (Jones et al., 1998) or GP-UCB (Srinivas et al., 2010) is used as an acquisition function;
@@
Note that most configurations for the baselines follow the configurations described in the work (Song et al., 2022).
@@
5.1 A Scenario with Unlabeled Point Sampling
Synthetic Benchmarks.
@@ 5.2 Scenarios with Fixed-Size Pools
Comparisons of our methods and the existing methods are carried out in these hyperparameter optimization benchmarks (Klein & Hutter, 2019), as reported in Figure 5. We can benchmark a variety of machine learning models, which are defined with specific hyperparameters and trained on one of four datasets: naval propulsion, protein structure, Parkinson’s telemonitoring, and slice localization. There exist 62,208 models, which are used as a pool in this paper, for each dataset. Our algorithms show superior performance compared to other approaches. Similar to the synthetic functions, we argue that the adoption of a predefined pool leverages its performance. In some cases, Gaussian process-based Bayesian optimization is better than our methods.
@@
Missing details of all the experiments are described in Section D.
@@ 6 Discussion
We choose a distribution for unlabeled point sampling as the truncated multivariate normal distribution in order to satisfy the cluster assumption. To analyze our algorithm thoroughly, we demonstrate the effects of the number of sampled points and sampling distributions for unlabeled points in Section F, varying the number of unlabeled points and utilizing diverse sampling distributions, i.e., uniform distributions, Halton sequences (Halton, 1960), and Sobol’ sequences (Sobol’, 1967).
@@
Because the computational complexity of label propagation and label spreading depends on a pool size, we need to reduce a pool size appropriately in order to speed up the algorithms. Therefore, we uniformly sample a subset of the fixed-size set, which is used as unlabeled points. Figure 14 reports elapsed times over subset sizes for pool sampling. Larger subsets make the framework slower as expected. More detailed analysis on the effects of pool sampling is demonstrated in Section G.
@@
We study the effects of a threshold ratio $\zeta$ in order to understand how we can choose $\zeta$. As shown in Figure 16, both small $\zeta$ and large $\zeta$ generally represent worse performance; see the results with $\zeta=0.01$ and $\zeta=0.8$. While it has to depend on optimization problems, $\zeta=0.33$ and $\zeta=0.5$ are generally appropriate choices according to this analysis. The details of this study can be found in Section H and Figure 16.

### v5

@@ 1 Introduction
To make use of semi-supervised classifiers, we take into account two distinct scenarios with unlabeled point sampling and with a predefined fixed-size pool. For the first scenario, we randomly sample unlabeled data points from the truncated multivariate normal distribution using a minimax tilting method (Botev, 2017), to allow for the possibility of adopting a cluster assumption (Seeger, 2000). Finally, we demonstrate that our method shows superior performance compared to the exiting methods in diverse experiments including synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and 64D minimum multi-digit MNIST search. Note that Tabular Benchmarks, NATS-Bench, and 64D minimum multi-digit MNIST search have access to fixed-size pools. To validate our methods, we provide thorough analyses on the components of DRE-BO-SSL in Section 6 and the appendices.
@@ 1.1 Over-Exploitation Problem
By the aforementioned definition, at early iterations of Bayesian optimization, a supervised classifier tends to overfit to a small size of $\mathcal{D}_{t}$ due to a relatively large model capacity. This consequence makes a Bayesian optimization algorithm highly focus on exploitation. Similar to our observation, the imbalance of exploration and exploitation in the DRE-based approaches is also discussed in the recent work by Oliveira et al. (2022); Pan et al. (2024). Moreover, the consequence mentioned above is different from the characteristics of probabilistic regression-based Bayesian optimization because the regression-based methods are capable of exploring unseen regions by dealing with uncertainties. Besides, even though a threshold ratio $\zeta$ might be able to mitigate this problem, an overconfident supervised classifier is likely to keep getting stuck in a local optimum as $y^{\dagger}$ cannot change dramatically; more detailed analysis on $\zeta$ is provided in Sections 6 and I.
@@ 2 Background and Related Work
[…]
@@
Whereas knowing a data distribution $p(\mathbf{x})$ is important, it is difficult to directly estimate $p(\mathbf{x})$ (Sugiyama et al., 2012). For specific machine learning problems such as importance sampling (Kloek & van Dijk, 1978) and mutual information estimation (Bishop, 2006), we can bypass direct density estimation and estimate a density ratio. More recently, Rhodes et al. (2020) tackle a density-chasm problem using telescopic density-ratio estimation, which replaces an original problem with a set of logistic regression problems. Since the work by Rhodes et al. (2020) suffers from the issue on distribution shift, the recent work by Srivastava et al. (2023) proposes a density ratio estimation method with multinomial logistic regression, which is capable of mitigating the distribution shift using multi-class classification. In Bayesian optimization, Bergstra et al. (2011) have proposed a strategy with tree-structured Parzen estimators to estimate a density ratio as an alternative to probabilistic regression-based acquisition functions. In addition, the existing work by Tiao et al. (2021); Song et al. (2022) suggests methods with binary classifiers in order to estimate class probabilities as a density ratio; see Section 3 for the details of this literature.
@@ 3 DRE-based Bayesian Optimization
where $\zeta=p(y\leq y^{\dagger})\in[0,1)$ is a threshold ratio. We need to find a maximizer of (1), by optimizing the following composite function: $h(p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})/p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t}))$, where $h(x)=(\zeta+(1-\zeta)x^{-1})^{-1}$. Since $h$ is a strictly increasing function, we can directly maximize $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})/p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t})$. In the previous work by Bergstra et al. (2011), two tree-structured Parzen estimators are used to estimate the respective densities, $p(\mathbf{x}\mid y\leq y^{\dagger},\mathcal{D}_{t})$ and $p(\mathbf{x}\mid y>y^{\dagger},\mathcal{D}_{t})$.
While the work by Bergstra et al. (2011) utilizes two distinct tree-structured Parzen estimators, Tiao et al. (2021) propose a method to directly estimate (1) using class-probability estimation (Qin, 1998; Sugiyama et al., 2012), which is called BORE. Since it can be formulated as a problem of binary classification in which Class 1 is a group of the top $\zeta$ of $\mathcal{D}_{t}$ and Class 0 is a group of the bottom $1-\zeta$ of $\mathcal{D}_{t}$ in terms of function evaluations, the acquisition function defined in (1) induces the following:
@@
Therefore, a class probability over $\mathbf{x}$ for Class 1 is considered as an acquisition function; it is simply derived by Tiao et al. (2021) as the following:
@@
which is versatile for any non-negative utility function $u(y;y^{\dagger})$, where $f$ is a strictly convex function, $f^{\prime}$ is the derivative of $f$, and $f^{*}$ is the convex conjugate of $f$. By the properties of LFBO, it is equivalent to an expected utility-based acquisition function. Along with the general framework, Song et al. (2022) prove that BORE is equivalent to the probability of improvement (Kushner, 1964) and LFBO with $u(y;y^{\dagger})=\mathbb{I}(y\leq y^{\dagger})$ is also equivalent to the probability of improvement, where $\mathbb{I}$ is an indicator function. Moreover, they show that their method with $u(y;y^{\dagger})=\max(y^{\dagger}-y,0)$, which can be implemented as weighted binary classification, is equivalent to the expected improvement (Jones et al., 1998).
@@ 4 DRE-based Bayesian Optimization with Semi-Supervised Learning
We adopt a multi-started local optimization technique, e.g., L-BFGS-B (Byrd et al., 1995), to solve (6). However, a flat landscape of $\pi_{\widehat{\mathbf{C}}_{t}}(\mathbf{x};\zeta,\mathcal{D}_{t},\mathbf{X}_{u})$ over $\mathbf{x}$ may occur, so that optimization performance can be degraded. To tackle this issue, a simple heuristic of randomly selecting a query point among points with identical highest class probabilities complements our method. Since the multi-started technique is utilized and the output of $\pi_{\widehat{\mathbf{C}}_{t}}$ is bounded in [0,1][0,1][ 0 , 1 ], a flat landscape is easily recognized by comparing the outcomes of the multi-started strategy.
@@ 4.1 Label Propagation and Label Spreading
where $\beta$ is a free parameter given. As discussed in the work by Zhu & Ghahramani (2002), we can learn $\beta$ in (8) by minimizing an entropy for propagated labels $\widehat{\mathbf{C}}$:
@@
See Figure 4 for the results of learning $\beta$ and Section F for analysis on the effects of $\beta$. By (8), we compute a transition probability from $\mathbf{x}_{j}$ to $\mathbf{x}_{i}$ by $p_{ij}=w_{ij}/\sum_{k=1}^{n_{l}+n_{u}}w_{kj}$. Note that similarities $\mathbf{W}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ and transition probabilities $\mathbf{P}\in\mathbb{R}^{(n_{l}+n_{u})\times(n_{l}+n_{u})}$ are defined, where $[\mathbf{W}]_{ij}=w_{ij}$ and $[\mathbf{P}]_{ij}=p_{ij}$. Moreover, by the definition of $p_{ij}$, $\mathbf{P}=\mathbf{D}^{-1}\mathbf{W}$, where $\mathbf{D}$ is a degree matrix whose diagonal entry is defined as the following:
@@
for $i\in[2]$, where $\mathbf{w}\in\mathbb{R}^{n_{l}+n_{u}}$ is similarities of $\mathbf{x}$ and $\mathbf{X}$ by (8). $[\widehat{\mathbf{C}}]_{:i}$ denotes $i$th column of $\widehat{\mathbf{C}}$. This inductive model is better than or equivalent to other classifiers without unlabeled data in certain circumstances; its analysis is provided in Section D.
@@ 4.2 Unlabeled Point Sampling
As described above, if unlabeled points are not available, we need to generate unlabeled points under a transductive learning scheme. However, it is a challenging problem unless we know a landscape of $\pi_{\mathbf{C}}$ adequately. Many studies by Seeger (2000); Rigollet (2007); Singh et al. (2008); Ben-David et al. (2008); Carmon et al. (2019); Wei et al. (2020); Zhang et al. (2022) investigate how unlabeled data can affect a model and whether unlabeled data helps improve the model or not.
@@
Assumption 4.1 (Cluster assumption in the work by Seeger (2000)).
@@
where $\mathbf{l},\mathbf{u}\in\mathbb{R}^{d}$ are lower and upper bounds, $\boldsymbol{\Sigma}=\mathbf{A}\mathbf{A}^{\top}$ is a covariance matrix, $\mathbb{I}$ is an indicator function, and $Z\sim\mathcal{N}(\boldsymbol{0},\mathbf{I}_{d})$ is a random variable. It is challenging to calculate a denominator of (15), $P(\mathbf{l}\leq\mathbf{A}^{\top}Z\leq\mathbf{u})$, and simulate from $f(\mathbf{z})$ because an integration of the denominator and an accept-reject sampling strategy from $f(\mathbf{z})$ are cumbersome in this multi-dimensional case. To effectively sample from the truncated multivariate normal distribution, we adopt the minimax tilting method (Botev, 2017). Compared to the method by Genz (1992), it yields a high acceptance rate and accurate sampling. In this paper $\boldsymbol{\Sigma}$ is set as an identity matrix, and $\mathbf{l}$ and $\mathbf{u}$ are determined by a search space. We will provide more detailed discussion on point sampling in Sections 6 and G.
@@ 5 Experiments
We compare baseline methods with DRE-BO-SSL in the following optimization problems: synthetic benchmarks for a scenario with unlabeled point sampling, and synthetic benchmarks, Tabular Benchmarks (Klein & Hutter, 2019), NATS-Bench (Dong et al., 2021), and minimum multi-digit MNIST search for a scenario with a fixed-size pool. Note that Tabular Benchmarks, NATS-Bench, and minimum multi-digit MNIST search are defined with a fixed number of possible solution candidates, which implies that they are considered as combinatorial optimization problems. By following the previous work by Tiao et al. (2021); Song et al. (2022), we set a threshold ratio as $\zeta=0.33$ for all experiments; the effects of a threshold ratio $\zeta$ are analyzed in Sections 6 and I. To solve (6), we use L-BFGS-B (Byrd et al., 1995) with 1,000 different initializations. All experiments are repeated 20 times with 20 fixed random seeds, where 5 initial points are given to each experiment. The sample mean and the standard error of the sample mean are reported. Other missing details including the details of the competitors of our methods are presented in Section E.
@@
  • Gaussian process, EI and UCB: It is a Bayesian optimization strategy, which is defined with Gaussian process regression with the Matérn 5/2 kernel (Rasmussen & Williams, 2006), where expected improvement (Jones et al., 1998) or Gaussian process upper confidence bound (Srinivas et al., 2010) is used as an acquisition function;
@@
Note that most configurations for the baselines follow the configurations described in the work by Song et al. (2022).
@@
5.1 A Scenario with Unlabeled Point Sampling
Synthetic Benchmarks.
@@ 5.2 Scenarios with Fixed-Size Pools
Comparisons of our methods and the existing methods are carried out in these hyperparameter optimization benchmarks (Klein & Hutter, 2019), as in Figure 5. We can benchmark a variety of machine learning models, which are defined with specific hyperparameters and trained on one of four datasets: naval propulsion, protein structure, Parkinson’s telemonitoring, and slice localization. There exist 62,208 models, which are used as a pool in this paper, for each dataset. Our algorithms show superior performance compared to other approaches. Similar to the synthetic functions, we argue that the adoption of a predefined pool leverages its performance. In some cases, the Gaussian process-based strategy is better than our methods.
@@
Missing details of these experiments are shown in Section E.
@@ 6 Discussion
We choose a distribution for unlabeled point sampling as the truncated multivariate normal distribution in order to satisfy the cluster assumption. To analyze our algorithm thoroughly, we demonstrate the effects of the number of sampled points and sampling distributions for unlabeled points in Section G, varying the number of unlabeled points and utilizing diverse sampling distributions, i.e., uniform distributions, Halton sequences (Halton, 1960), and Sobol’ sequences (Sobol’, 1967).
@@
Because the computational complexity of label propagation and label spreading depends on a pool size, we need to reduce a pool size appropriately in order to speed up the algorithms. Therefore, we uniformly sample a subset of the fixed-size set, which is used as unlabeled points. Figure 14 reports elapsed times over subset sizes for pool sampling. Larger subsets make the framework slower as expected. More detailed analysis on the effects of pool sampling is demonstrated in Section H.
@@
We study the effects of a threshold ratio $\zeta$ in order to understand how we can choose $\zeta$. As shown in Figure 16, both small $\zeta$ and large $\zeta$ generally represent worse performance; see the results with $\zeta=0.01$ and $\zeta=0.8$. While it has to depend on optimization problems, $\zeta=0.33$ and $\zeta=0.5$ are generally appropriate choices according to this analysis. The details of this study can be found in Section I and Figure 16.